# Conditional Text-to-Image Generation with Reference Guidance

### v1

@@ 4.5 Ablation Study
In this section, we present ablation studies to gain deeper insights into the proposed method. For additional additional ablation please refer to Appendix B.
@@ 4.6 Quantitative Results
We provide comparison results with SD [45], ControlNet [65], DeepFloyd [2], and TextDiffuser [12]. For ControlNet, we use the model trained to be conditioned on the Canny Edge map, and we provide Canny Edge maps of rendered text images or logo images as input conditions during inference. We present results for our method both with and without the scheduled CFG guidance described in Section 3.4.
English Images. For the English image generation task, our method achieves the best performance on all the OCR-related metrics, including OCR Accuracy and F-1 Score (Table 4). Additionally, our method also achieves the best CLIP score, demonstrating that our model is not only capable of generating accurate images but also best aligns with the text prompts. Besides, our method achieves comparable results of FID with DeepFloyd which achieves lower accuracy than ours by a large margin. When the proposed CFG scheduling scheme is applied, we obtain further improvement in accuracy, with a decrease in FID performance. We provide additional analysis on CFG scheduling at Appendix B.
@@
[…]
SD 51.295 0.3015 0.03 2.14
ControlNet 51.485 0.3424 23.90 58.65
DeepFloyd 34.902 0.3267 2.62 17.62
TextDiffuser 38.758 0.3436 56.09 78.24
Ours 38.593 0.3454 58.26 79.15
Ours+ 42.192 0.3434 61.73 80.08
Table 4: Comparison results of English generation on MARIO-Eval benchmark. Acc and F-1 denote OCR Accuracy and F1-score, respectively. $+$ denotes the results with scheduled classifier guidance. Accuracy and F-1 are denoted in [%].
Method Latin Non-latin
[…]
SD 114.34 0.3032 0.38 1.87 113.07 0.3067 0.00 0.00
ControlNet 140.03 0.3010 10.63 43.03 136.54 0.3038 12.00 40.09
TextDiffuser 115.51 0.2922 58.25 68.73 130.44 0.2879 0.00 0.00
Ours 117.20 0.2952 56.38 76.68 116.79 0.3059 23.80 38.96
Ours+ 117.96 0.2803 64.56 79.09 117.09 0.2940 29.20 41.78
Table 5: Evaluation results of the MLT image generation. Latin languages: English, German, French, and Italian. Non-latin languages: Bengali, Hindi, Greek, Russian, and Thai. Result denoted by $+$ applies CFG scheduling with $\rho_{\text{speed}}=0.2$, $\gamma=3.5$ and $\omega_{\text{all}}=4$. Accuracy and F-1 are denoted in [%].
@@
Table 6: Evaluation results for logo image generation. Result denoted by $+$ applies CFG scheduling with $\rho_{\text{speed}}=0.2$, $\gamma=3.5$ and $\omega_{\text{all}}=4$ for CFG scheduling. Accuracy and F-1 are denoted in [%].
@@ 4.7 Qualitative Results
We present side-by-side results in Figure 6 for direct comparisons of the image quality of different methods. Our method achieves both faithful blending of target subjects to the scene and strong correspondence to the text prompt in all experiments. We also provide logo image generation results when unseen logo images are given (Figure 7), which further validates the effectiveness of the proposed method and the model’s ability to generalize across novel instances. Moreover, our method is highly flexible and can be easily extended to text editing tasks (Figure 8). Following [12], we augment the channel of input latent by concatenating encoded latent of the masked image. We provide more details and additional qualitative results in Appendix C.
@@
Figure 8: Text inpainting results. Regions denoted in yellow are masked during the inference.
@@ Appendix A Additional Technical Details
A.1 Low-Rank Adaptation Details
We elaborate on the details of Low-Rank Adaptation (LoRA). Instead of fine-tuning the entire model, we attempt to fine-tune the residual of the model parameters for efficiency:
$\theta_{l}^{\prime}=\theta_{l}+\alpha\Delta\theta_{l},$ (1)
where $\alpha$ denotes the merging ratio, and $\theta_{l}\in\mathbb{R}^{d_{l_{1}}\times{d_{l_{2}}}}$ denotes $l^{\text{th}}$ layer of the model. We set $\alpha=1$. Here, the $\Delta\theta_{l}$ can be further decomposed into low-rank matrices as follows:
$\Delta\theta_{l}=A_{l}B_{l}^{T}.$ (2)
where $A_{l}\in\mathbf{R}^{d_{l_{1}}\times r}$ and $B_{l}\in\mathbf{R}^{d_{l_{2}}\times r}$, and $r$ is a small value compared to $l_{d_{1}}$, and $l_{d_{2}}$. We fine-tune the low-rank matrices for efficiency, and this learning scheme is adopted for all the layers of the diffusion model, including fully connected layers, convolutional layers, and cross-attention layers. We study the impact of the size of $r$ in Table G of the main paper.
A.2 Denoised Input Reconstruction
The denoised latent features are reconstructed using the following formula [21]:
$\hat{\mathbf{z}}^{0}=(\mathbf{z}^{t})-\sqrt{1-\bar{\alpha_{t}}}\epsilon_{\theta}(\mathbf{z}^{t}))/\sqrt{\bar{\alpha}_{t}},$ (3)
where $\hat{\mathbf{z}}^{0}$ is the reconstructed latent, $\mathbf{z}^{t}$ is the corrupted latent at timestep $t$, $\epsilon_{\theta}$ is the predicted epsilon, $t$ is the timestep, $\bar{\alpha_{t}}=\prod_{s=1}^{t}\alpha_{s}$, $\alpha_{t}=1-\beta_{t}$, and $\beta_{t}$ is the forward process variance, respectively.
A.3 Auxiliary Network Architecture
We provide details of the auxiliary recognition network architecture. The auxiliary network is composed of the encoder and the recognition head. The encoder for the text image generation and logo image generation share the same encoder structure. The shared encoder network consists of multiple encoding blocks, each sharing the structure of the blocks in the VAE network of the diffusion model. More specifically, each encoder block comprises a GroupNorm layer, two Residual blocks, and another GroupNorm layer, where the number of groups is set to 16. We use four encoding blocks for both text recognition and log recognition networks. Except for the last encoder block, the spatial dimension is down-sampled by the factor of 2 using the convolution layer with stride=2.
Text recognition head. The text recognition head is composed of four cross-attention layers, where the cross-attention is computed between the sequence of input image tokens and the sequence of positional embeddings corresponding to the characters of the output text word. The cross-attention we compute here can be formulated as:
[…]
@@ A.3 Auxiliary Network Architecture
Logo recognition head. The recognition head for the auxiliary logo recognition network is composed of a single fully connected layer, where the output dimension corresponds to the size of the logo set.
A.4 Additional Details of Augmented CFG
Classifier-Free Guidance [20] considers sharpened posterior distribution $P_{\theta}(\mathbf{z}|\textbf{c})\propto P_{\theta}(\mathbf{z})P_{\theta}(\textbf{c}|\mathbf{z})^{\omega}$. Using Bayes rule for some timestep $t$,
$\begin{split}\nabla_{\mathbf{z}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{c})&=\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\\ &+\omega(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{c})-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})).\end{split}$ (5)
@@ A.4 Additional Details of Augmented CFG
$\hat{\epsilon_{\theta}}^{t}:=\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)+\omega(\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)-\epsilon_{\theta}(\mathbf{z}^{t},\varnothing,t)),$ (6)
@@
Similarly, we model the posterior distribution that can be represented as,
$\begin{split}P_{\theta}(\mathbf{z}|\textbf{s},\textbf{m},\textbf{c})&\propto P_{\theta}(\mathbf{z})P_{\theta}(\textbf{s},\textbf{m},\textbf{c}|\mathbf{z})^{\omega_{\text{all}}}\\ &\quad\cdot P_{\theta}(\mathbf{z})P_{\theta}(\textbf{s},\textbf{m}|\mathbf{z})^{\omega_{\text{prompt}}}P_{\theta}(\textbf{c}|\mathbf{z})^{\omega_{\text{ref}}}.\end{split}$ (7)
@@
$\begin{split}\nabla_{\mathbf{z}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{s},\textbf{m},\textbf{c})&=\nabla_{\mathbf{z}_{t}}\log P(\mathbf{z}_{t})\\ &\quad+\omega_{\text{all}}\left(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{s},\textbf{m},\textbf{c})-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\right)\\ &\quad+\omega_{\text{ref}}\left(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{s},\textbf{m},\varnothing)-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\right)\\ &\quad+\omega_{\text{prompt}}\left(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\varnothing,\varnothing,\textbf{c})-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\right).\end{split}$ (8)
@@
$\begin{split}\hat{\epsilon}^{t}&=\underbrace{\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)}_{\text{Unconditional}}\\ &\underbrace{+\omega_{\text{prompt}}\Big{[}\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\mathbf{c},t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{Caption Guidance}}\\ &+\underbrace{\omega_{\text{ref}}\Big{[}\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\varnothing,t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{Reference Guidance}}\\ &+\underbrace{\omega_{\text{all}}\Big{[}\epsilon(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\mathbf{c},t)-\epsilon(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big{]}}_{\text{All Guidance}},\end{split}$ (9)
A.5 Additional Details of Datasets
@@
MARIO-10M [12] is a compilation of data from various publicly accessible sources, including the LAION-400M [50], TMDB [6], and Open Library datasets. The images within this dataset are filtered by a text detector and annotated by an OCR model. A training set containing 10 million images with English-only annotations are utilized for training.
@@
ICDAR2019 [14] stands as a multi-lingual text (MLT) dataset initially curated for text detection and recognition. This dataset contains 10,000 training images and provides coordinates of each text word with corresponding character labels. We use a subset with six languages, including English, French, German, and Italian to train our MLT image generation from this dataset.
@@
FlickrLogos-32 [46] contains 8,240 images with 32 logo brands. This dataset was originally collected for the logo retrieval and logo detection/recognition tasks. We use a subset of 2,240 images that contain at least one logo as the training set for the logo image generation task.
@@
Logos in the wild (LITW) [57] comprises 11,054 images obtained from the Google image search engine, featuring 871 unique brands. We filter the dataset by excluding images with logos smaller than 45 pixels high and images with no logo. After filtering, a total of 4,206 images are obtained for training.
@@
LAION-Aesthetics [4]. LAION-5B [51] is a large-scale dataset of image-caption pairs. LAION-Aesthetics is the subset of LAION-5B, filtered by a trained aesthetic score predictor, and contains only high aesthetic score (above 6.5) images. This dataset does not provide any annotations of text or logo, and is included in our training to prevent overfitting and catastrophic forgetting.
@@
Aesthetic text dataset (AText). We collect an additional dataset for English text generation with high-aesthetic and high-artistic images from the Internet [5]. We measure the aesthetic scores of the images using a pretrained model from [51] and exclude those with low aesthetic scores. We run an OCR model to detect text regions and annotate the text within the regions. Example images are shown in Figure 5.
@@
Synthetic logo/text dataset. Due to the limited numbers of MLT and logo training samples from public datasets, we augment the training set with synthesis. We use high-resolution images from public datasets as the background and render the target text or logo. For the multi-lingual experiment, we use an E-book text corpus [1] and translate the text into the target languages of interest. For the logo image experiment, we use publicly available icon images [3] as the target objects to be rendered on the background images (Figure 4). In total, we obtain 23,892 synthetic logo images and 184,912 synthetic MLT images.
A.6 Text Image Editing Details

### v2

@@ Abstract
1004.1 Datasets
We explain the dataset that we use for training. Additional details are available in Sections A.3 and A.4.
English. For training English plug-in models, we first train the model with MARIO-10M [textdiffuser] and further fine-tune the model by including samples from LAION-Aesthetic [laion] and additional aesthetic text images (AText) that we collect (Figure 5). We use MARIO-Eval [textdiffuser] benchmark for evaluation.
Multi-Lingual Text. For the MLT plug-in, we finetune the English plug-in model trained with MARIO-10M using the merged set of real images from ICDAR2019 [icdar2019], and synthetic images that we generated. Due to the lack of pre-existing benchmarks for MLT image generation, we developed a new benchmark for model evaluation on nine different languages, having 2,500 prompts in total. Details of the benchmark are available in Section A.4.
Logos. Similar to the MLT plug-in, we also finetune the English plug-in model trained with MARIO-10M using the merged set of real images from the merged set of FlickrLogos-32 [flickr27] and Logos in the wild (LITW) [litw], and the synthetic images that we generated. For evaluation
1004.2 Evaluation Metrics
We use Fréchet inception distance (FID) to assess the image quality, and we evaluate the semantic alignment by computing the CLIP Score between the text prompts and the generated images and measuring the aesthetic scores of the generated images obtained from a pretrained model [laion].
Text images. To measure the accuracy of the generated text, we use Accuracy and F1-score. We apply different OCR Engines for English generation tasks and MLT image generation tasks. We use the Microsoft Read API for English, and the Google Cloud Vision API for MLT image generation tasks. We consider a detected text correct if it is exactly matched with a keyword, and the F-1 score is defined as the harmonic mean of precision and recall.
Logo images. To measure the accuracy of the logo image generation, we train a Mask-RCNN [mask_rcnn] on the merged set of FlickrLogos-32 [flickr32] and LITW dataset [litw]. We report the Accuracy and F1-score of the logo detection model applied to an image set generated by models.
[…]
SD 51.29 0.3015 0.03 2.14
ControlNet 51.48 0.3424 23.90 58.65
DeepFloyd 34.90 0.3267 2.62 17.62
TextDiffuser 38.75 0.3436 56.09 78.24
SD3 37.21 0.3424 9.62 49.55
SDXL 58.54 0.3242 2.16 20.33
Flux 41.25 0.3198 54.98 52.35
GlyphControl 61.10 0.3411 27.04 64.00
AnyText 52.39 0.3426 18.03 60.74
Ours 38.59 0.3454 58.26 79.15
Ours+ 42.19 0.3434 61.73 80.08
Table 4: Comparison results of English generation on MARIO-Eval benchmark. Acc and F-1 denote OCR Accuracy and F1-score, respectively. ++ denotes the results with scheduled classifier guidance. Accuracy and F-1 are denoted in [%].
1004.3 Implementation Details.
We first fine-tune the pre-trained Stable Diffusion (SD) V2.1 [SD] on the MARIO-10M dataset [textdiffuser]. For English image generation, we further train the model using the AText dataset we collect. For MLT and Logo image generation, we train the model with merged sets of synthetic images and real images. We set batch size=10 for each of the 8 GPUs. We use AdamW [adamw] with a learning rate of 1e-4 to tune the parameter. For MLT image generation, we use a character set containing 847 characters from the nine languages. We use a rank of 32 for all LoRA. We use 100 steps of DDIM [ddim] sampling. For all the experiments, we set the recognition loss weight $\lambda$ as 0.025.
Method Latin Non-latin
[…]
SD 114.34 0.3032 0.38 1.87 113.07 0.3067 0.00 0.00
GlyphControl 115.51 0.2922 58.25 68.73 130.44 0.2879 0.00 0.00
AnyText 115.51 0.3100 10.38 48.87 130.44 0.3155 4.50 17.42
SD3 134.31 0.2983 25.06 46.13 133.24 0.3090 0.11 0.11
SDXL 126.49 0.3041 16.18 13.90 132.13 0.3101 0.00 0.00
ControlNet 140.03 0.3010 10.63 43.03 136.54 0.3038 12.00 40.09
Ours 117.20 0.2952 56.38 76.68 116.79 0.3059 23.80 38.96
Ours+ 117.96 0.2803 64.56 79.09 117.09 0.2940 29.20 41.78
Table 5: Evaluation results of the MLT image generation. Latin languages: English, German, French, and Italian. Non-Latin languages: Bengali, Hindi, Greek, Russian, and Thai. Result denoted by ++ applies CFG scheduling with $\rho_{\text{speed}}=0.2$, $\gamma=3.5$ and $\omega_{\text{all}}=4$. Accuracy and F-1 are denoted in [%].
@@
In this section, we present ablation studies to gain deeper insights into the proposed method. We provide additional ablation studies in Appendix B.
@@
We provide comparison results with SD [SD], SD3 [sd3], SDXL [podell2023sdxl], Flux [flux], GlyphControl [glyphcontrol], AnyText [anytext], ControlNet [controlnet], DeepFloyd [deepfloyd], and TextDiffuser [textdiffuser]. For ControlNet, we use the model trained to be conditioned on the Canny Edge map, and we provide Canny Edge maps of rendered text images or logo images as input conditions during inference. For log generation, we also compare with BLIP-Diffusion [blipdiff], IP-Adapter [ipadapter], and MS-Diffusion [msdiff]. We present results for our method both with and without the scheduled CFG guidance described in Section 1003.4.
English Images. For the English image generation task, our method achieves the best performance on all the OCR-related metrics, including OCR Accuracy and F-1 Score (Table 4). Additionally, our method also achieves the best CLIP score, demonstrating that our model is not only capable of generating accurate images but also best aligns with the text prompts. Besides, our method achieves comparable results of FID with DeepFloyd, which achieves lower accuracy than ours by a large margin. When the proposed CFG scheduling scheme is applied, we obtain further improvement in accuracy, with a decrease in FID performance. We provide additional analysis on CFG scheduling at Appendix B.
@@
Table 6: Evaluation results for logo image generation. Result denoted by ++ applies CFG scheduling with $\rho_{\text{speed}}=0.2$, $\gamma=3.5$ and $\omega_{\text{all}}=4$ for CFG scheduling. Accuracy and F-1 are denoted in [%].
@@
We present side-by-side results in Figure 6 for direct comparisons of the image quality of different methods. Our method achieves both faithful blending of target subjects to the scene and strong correspondence to the text prompt in all experiments. We also provide logo image generation results when unseen logo images are given (Figure 7), which further validates the effectiveness of the proposed method and the model’s ability to generalize across novel instances. Moreover, our method is highly flexible and can be easily extended to text editing tasks (Figure 8). Following [textdiffuser], we augment the channel of the input latent by concatenating the encoded latent of the masked image. We provide more details and additional qualitative results in Appendix C.
@@
Figure 8: Text editing results. Regions denoted in yellow are masked and edited.
@@ Appendix A Additional Technical Details
A.1 Auxiliary Network Architecture
We provide details of the auxiliary recognition network architecture. The auxiliary network is composed of the encoder and the recognition head. The encoders for the text image generation and logo image generation share the same encoder structure. The shared encoder network consists of multiple encoding blocks, each sharing the structure of the blocks in the VAE network of the diffusion model. More specifically, each encoder block comprises a GroupNorm layer, two Residual blocks, and another GroupNorm layer, where the number of groups is set to 16. We use four encoding blocks for both text recognition and log recognition networks. Except for the last encoder block, the spatial dimension is down-sampled by the factor of 2 using the convolution layer with stride=2.
Text recognition head.
The text recognition head is composed of four cross-attention layers, where the cross-attention is computed between the sequence of input image tokens and the sequence of positional embeddings corresponding to the characters of the output text word. The cross-attention we compute here can be formulated as:
$\text{Attention}(\mathbf{r},\tilde{\mathbf{z}},\tilde{\mathbf{z}})=\text{ softmax}(\frac{\mathbf{r}\tilde{\mathbf{z}}^{T}}{\sqrt{d_{k}}})\tilde{\mathbf{z}}$, (1)
@@ A.1 Auxiliary Network Architecture
Logo recognition head.
The recognition head for the auxiliary logo recognition network is composed of a single fully connected layer, where the output dimension corresponds to the size of the logo set.
Text Recognition Loss.
The text recognition loss that we use is defined as below,
$\small\textbf{r}_{k}$ $=\text{RoIAlign}(\hat{\mathbf{z}}^{0},\mathbf{B}_{k})$ (2)
$\textbf{o}_{k}$ $=\psi_{\text{recog}}(\textbf{r}_{k})$ (3)
$\mathcal{L}_{\text{recog}}$ $=\frac{-1}{L}\sum_{j=1}^{L}y_{j}\log\textbf{o}_{k,j},$ (4)
where $\textbf{B}_{k}$ is the bounding box label for $k$-th region in the image, $\textbf{o}_{k}$ is the network output, and $L$ denotes the length of the word in the $k$-th region. $j$ is an enumerator for each character in the word. We predict the probability distribution over the character set at each character position $j$ denoted as $\textbf{o}_{k},j\in\mathbb{R}^{|C|}$. For each position, we apply cross-entropy loss between the one-hot label $y_{j}\in\mathbb{R}^{|C|}$. For MLT loss, we use the same loss design, but the output dimension is extended to 847 to cover all the characters of the target languages.
A.2 Additional Details of Augmented CFG
Classifier-free guidance [cfg]
considers sharpened posterior distribution $P_{\theta}(\mathbf{z}|\textbf{c})\propto P_{\theta}(\mathbf{z})P_{\theta}(\textbf{c}|\mathbf{z})^{\omega}$. Using Bayes’ rule for some timestep $t$,
$\small\begin{split}\nabla_{\mathbf{z}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{c})&=\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\\ &+\omega(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{c})-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})).\end{split}$ (5)
@@ A.2 Additional Details of Augmented CFG
$\small\hat{\epsilon_{\theta}}^{t}:=\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)+\omega(\epsilon_{\theta}(\mathbf{z}^{t},\mathbf{c},t)-\epsilon_{\theta}(\mathbf{z}^{t},\varnothing,t)),$ (6)
@@
Similarly we model the posterior distribution that can be represented as,
$\small\begin{split}P_{\theta}(\mathbf{z}|\textbf{s},\textbf{m},\textbf{c})&\propto P_{\theta}(\mathbf{z})P_{\theta}(\textbf{s},\textbf{m},\textbf{c}|\mathbf{z})^{\omega_{\text{all}}}\\ &\qquad\cdot P_{\theta}(\mathbf{z})P_{\theta}(\textbf{s},\textbf{m}|\mathbf{z})^{\omega_{\text{prompt}}}P_{\theta}(\textbf{c}|\mathbf{z})^{\omega_{\text{ref}}}.\end{split}$ (7)
@@
$\footnotesize\begin{split}\nabla_{\mathbf{z}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{s},\textbf{m},\textbf{c})&=\nabla_{\mathbf{z}_{t}}\log P(\mathbf{z}_{t})\\ &\qquad+\omega_{\text{all}}\left(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{s},\textbf{m},\textbf{c})-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\right)\\ &\qquad+\omega_{\text{ref}}\left(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\textbf{s},\textbf{m},\varnothing)-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\right)\\ &\qquad+\omega_{\text{prompt}}\left(\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t}|\varnothing,\varnothing,\textbf{c})-\nabla_{\mathbf{z}_{t}}\log P_{\theta}(\mathbf{z}_{t})\right).\end{split}$ (8)
@@
$\small\begin{split}\hat{\epsilon}^{t}&=\underbrace{\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)}_{\text{Unconditional}}\\ &\underbrace{+\omega_{\text{prompt}}\Big[\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\mathbf{c},t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]}_{\text{Caption Guidance}}\\ &+\underbrace{\omega_{\text{ref}}\Big[\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\varnothing,t)-\epsilon_{\theta}(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]}_{\text{Reference Guidance}}\\ &+\underbrace{\omega_{\text{all}}\Big[\epsilon(\text{concat}(\mathbf{z}^{t},\textbf{s},\textbf{m}),\mathbf{c},t)-\epsilon(\text{concat}(\mathbf{z}^{t},\varnothing,\varnothing),\varnothing,t)\Big]}_{\text{All Guidance}},\end{split}$ (9)
A.3 Additional Details of Datasets
MARIO-10M [textdiffuser]
@@ A.3 Additional Details of Datasets
is a compilation of data from various publicly accessible sources, including the LAION-400M [laion400m], TMDB [tmdb], and Open Library datasets. The images within this dataset are filtered by a text detector and annotated by an OCR model. A training set containing 10 million images with English-only annotations are utilized for training.
ICDAR2019 [icdar2019]
@@
stands as a multi-lingual text (MLT) dataset initially curated for text detection and recognition. This dataset contains 10,000 training images and provides coordinates of each text word with corresponding character labels. We use a subset with six languages, including English, French, German, and Italian to train our MLT image generation from this dataset.
FlickLogos-32 [flickr32]
@@
contains 8,240 images with 32 logo brands. This dataset was originally collected for the logo retrieval and logo detection/recognition tasks. We use a subset of 2,240 images that contain at least one logo as the training set for the logo image generation task.
Logos in the wild (LITW) [litw]
@@
comprises 11,054 images obtained from the Google image search engine, featuring 871 unique brands. We filter the dataset by excluding images with logos smaller than 45 pixels high and images with no logo. After filtering, a total of 4,206 images are obtained for training.
LAION-Aesthetics [laion_aes].
@@
LAION-5B [laion] is a large-scale dataset of image-caption pairs. LAION-Aesthetics is the subset of LAION-5B, filtered by a trained aesthetic score predictor, and contains only high aesthetic score (above 6.5) images. This dataset does not provide any annotations of text or logo, and is included in our training to prevent overfitting and catastrophic forgetting.
Aesthetic text dataset (AText).
@@
We collect an additional dataset for English text generation with high-aesthetic and high-artistic images from the Internet [pexels]. We measure the aesthetic scores of the images using a pretrained model from [laion] and exclude those with low aesthetic scores. We run an OCR model to detect text regions and annotate the text within the regions. Example images are shown in Figure 5.
Synthetic logo/text dataset.
@@
Due to the limited numbers of MLT and logo training samples from public datasets, we augment the training set with synthesis. We use high-resolution images from public datasets as the background and render the target text or logo. For the multi-lingual experiment, we use an E-book text corpus [corpus] and translate the text into the target languages of interest. For the logo image experiment, we use publicly available icon images [icons] as the target objects to be rendered on the background images (Figure 4). In total, we obtain 23,892 synthetic logo images and 184,912 synthetic MLT images.
A.4 Additional Details of Benchmarks
Benchmark for English generation.
We adopt the MARIO-Eval [textdiffuser] benchmark to assess the model performance in generating English scene-text images. The MARIO-Eval benchmark contains 5,414 prompts. This benchmark is composed of six different subsets, each a subset of multiple benchmarks, including DrawBenchText [drawtext], DrawTextCreative [drawtext], ChineseDrawText [glyphdraw], and Mario-10M test set [textdiffuser]. Following the protocol [textdiffuser], we exclude images generated from ChineseDrawText and DrawBenchText, hence 5,000 images are used for measuring FID.
Benchmark for multi-lingual generation.
Due to the lack of pre-existing benchmarks for MLT image generation, we developed a new benchmark for model evaluation on nine different languages. We create 25 prompt templates, such as “A raccoon holding a paper saying words.”. We exclude the non-English target words from the prompts, as they cannot be processed by the text encoder. Each prompt is designed to include up to three words in the image. Except for English, we allocate the rest of eight languages a set of 200 prompts, and 400 for English, yielding 2,500 prompts in total. For computing FID, we construct a set of images by randomly sampling a subset of the ICDAR2019 dataset.
Benchmark for logo generation.
We additionally construct a benchmark for logo image generation. Specifically, we create 25 template prompts, e.g., “A building displaying a [KEYWORD] sign.”, where [KEYWORD] here is the name of the logo. We obtain 150 logos in total, which is the union of the class of FlickrLogos-32 dataset [flickr27] and LITW dataset, yielding 1,500 prompts in total. For computing FID, we construct an image set sampled from the merged set of FlickrLogos-32 [flickr32] and LITW [litw] dataset.
A.5 Text Image Editing Details